OpenStudy (anonymous):

find three consecutive integers 103

OpenStudy (anonymous):

Your question isn't really clear. Could you state it more clearly? What does the "193" mean?

OpenStudy (anonymous):

what are two consecutive integers that equal 193

OpenStudy (anonymous):

When added?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it two consecutive integers or two consecutive even integers

OpenStudy (anonymous):

Two consecutive integers that equal 193: We'll call our first integer n. The other one is n +1. Here is our equation: n+n+1=193 -1 -1 ----------- 2n = 192 Divide both sides by two: n = 96 96 = n, which is your first integer. Your second integer is n+1, which is 96+1 = 97. So, your two integers are 96 and 97.